OpenStudy (a1234):

Function value: csc theta = 4 Constraint: sin theta < 0 Find tan theta. Is it -((sqrt15)/15)?

OpenStudy (anonymous):

\[\sin ~ and~ \csc ~have~same~sign.\] \[\csc ^2 \theta-\cot ^2 \theta=1\] \[4^2-\cot ^2\theta=1\] \[\cot ^2\theta=16-1=15\] \[\tan ^2\theta=\frac{ 1 }{ 15 }\] \[\sin \theta <0,means ~\theta~lies ~\in~3rd~or~4th~quadrant.\] if in 3rd quadrant tan theta is positive If in 4th quadrant tan theta is negative.